40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Sun Jul 26 20:37:50 2020, aaron, Update, ,  
    Reply  Mon Jul 27 12:11:49 2020, aaron, Update, Lab Work, Simplifying W optic bench 
Message ID: 2543     Entry time: Mon Jul 27 12:11:49 2020     In reply to: 2542
Author: aaron 
Type: Update 
Category: Lab Work 
Subject: Simplifying W optic bench 

PSOMA hardware inventory

I located the materials for stage 1 PSOMA on the West optics table. I recorded what we have in the hardware inventory, and what we don't have is flagged for purchase. I start by cleaning up the electronics rack, removing anything I think is not in use.

Electronics rack

  • Both laser current drivers are D1500207
  • The E laser has a TED 200 C temp controller. The W laser is not plugged in to TEC, but there is a spare ITC 502 combination controller on the rack, and cryo cavs has a additional ITC 510.
  • There are three universal PDH servo boxes on the W rack: 1x D0901351 and 2x D1700219-v1. All of the boxes need lids and proper rack mount hardware.
    • Innards
      • Inside the *1351, the board matches the front panel DCC number (serial number 1437)
      • Both the *0219 have matching boards: D1700192-v2, D1700195-v2, D1799182-v2, D1700131-v2
    •  
  • Power supplies
    • The +- 15V power strip is currently unplugged. I don't see a low noise DC supply on this rack, but there is a Tenma 72-6615 on the ground. The closed-ring connector of the power strip is incompatible with this supply, so likely it was powered off of something else.
    • There's a NewFocus 1901 +- 15V supply on the W table itself. I think this should be located not on the table? The outputs are 0.1 A max to banana, plus 2x 0.3A max to a coax cable. The NewFocus was only supplying the PDs on the table. 
    • Cryo cavs has a dedicated rack for its power supplies (2x Sorensen DCS33-33E), separated from the laser control and PDH electronics. Is this an important choice, and should we also move our power supplies?
    • On the bottom of the rack, there's a Mech-Tronics NIM power supply. Seems out of use, only a PD power supply and readout board is plugged in, but this has no corresponding PD. 
    • I found on the elog that Chris moved the two Sorensens from the cryo cantilevers (W) rack to the cryoaux (central, rolling) rack after the supplies Johannes bought in 2017 died (fans not running). Do these usually die right after their 2 year warranty expires, or were we especially hard on them?
      • sidenote, great to see a potential solution to cryo lab timing woes, and that the overheated Sorensens were causing cominaux hard drive failures. Thanks Chris!
      • I've added 2x Sorensen supplies to our purchase list for PSOMA
    • For now, I moved all existing power supplies onto the electronics rack, and power the strip with the Tenma. In the future, we can discuss a remote location for power supplies, and will buy 2x additional Sorensens.
  • Misc
    • a bunch of old and open circuit boards, in cases compatible with the NIM supply. One power amp II; one PDH2 servo board (D1100996); 2 channel low noise amplifier (contains D1101396-v2 and another hand made board); a general filter board claiming to be from the gyro experiment. 
    • There were a few loose minicircuits components connected to the OCXO preamp (D1500064). These were providing 33.59 and 32.7 MHz modulation to Zach's experiments. I removed these from the preamp so we can give them a more permanent box later.

W optics bench

  • vacuum pump
    • I don't see the pumping station that used to be by the W optics bench. I don't remember using it elsewhere, and don't find any elog mentioning it being borrowed or moved. Has anyone seen this pump? It's one of the HiCube all-in-ones; it should have LIGO property tag C21832.
    • Similarly, what happened to the IR labs cryostat that housed the cantilever Q measurement (property tag)? I see maybe the glint of a cantilever in the short gryo vacuum can, and I have the cantilever Q HV driver and feedthrough, but no cryostat.
  • Photodiodes
    • PDA50B (Ge, 800-1800 nm)
    • 2x PDA255
    • Found 2 loose (unmounted) photodiodes wrapped in foil near the "optical contact razorblade test" setup (SE corner of W table).
    • NewFocus 1811, 1611
    • 1 home brewed QPD and breakout board, (PN on the circuit board is 10-00146 rev 1)
  • Put away or onto the workbench any loose optics, unused optomechanics, and most cables. Anything that wasn't a stock part I placed in a clear plastic box labeled 'Cryo Cantilevers' in the projects cabinet (2-part wooden cabinet on S wall). Stock parts went into the respective optics cabinets.

Misc

  • The lab temperature is still way too high. The temperature at the particle counter is reading 78, the thermostat still reads 71 F. The filter is not leaking substantially.j
  • Got the speaker working -- there was an aux adapter in the headphone port, so system was confused
  • Raymond requested a CF16 blank. I couldn't find one, but gave him a CF40-to-CF25 adapter that had a CF25 blank epoxied onto it, along with a CF16 to CF40 adapter and some gaskets. Here was the disinfection procedure for sending these items to QIL:
    1. Found a large plastic bag and disinfected inside and out with Phenomenal (TM) spray
    2. Disinfected the unpackaged UHV components with 70% isopropyl alcohol and a Kim wipe, wrapped them in foil, and placed in the disinfected bag
    3. Disinfect the outside of the bag with Phenomenal spray and leave outside the door to cryo for same-day pickup.
  • noted that pressure in cryo Q vacuum is 24 utorr after about a day valved off
  • Emailed Jordan and Liz about anemometer, to measure air flow in cryo. If we can compute the time for air to circulate through the cryo lab, we can reduce the length of time between lab entries (currently Shruti and I are leaving 1 day between our lab uses). 
  • Starting piling up seemingly unused components from W table onto the S workbench. These need a new home.
  • Found some historical documents, sticky notes from Zach. Took photos of these just in case (attached, later) and placed along with the miscellany on the workbench.
  • Checked with other W Bridge lab users before entering EE shop at 3:30pm. Disinfected surfaces I touched before and after use. I retrieved 3x 9V rechargeable batteries for the cryo lab multimeters.
  • Found a cardboard boxed fiber polarization controller (FPC032), which I placed in the NW optics cabinet sans box. Is this safe for storage (attachment 3)? What about things like Kim wipes, are those an acceptable amount of cardboard, or should they be eliminated?
  • Moved any cantilevers, etc into the silicon storage cabinet (NW corner cabinet)
ELOG V3.1.3-